# Dynamics-Aware Gaussian Splatting Streaming Towards Fast On-the-Fly Training for 4D Reconstruction

### v1

@@ 3.2 Selective Inheritance
In our pipeline, the shift and densification stages focus on the deformation of existing objects and the densification for emerging objects, respectively. Due to the strong temporal consistency inherent in the scene, the densified Gaussians from the previous frame, $\bm{\mathcal{G}}^{\textrm{densi}}_{t-1}$, are likely to persist and remain valid in the current frame, since they have been specifically optimized for the scene. Leveraging these frame-wise priors and dependencies for subsequent frames can significantly accelerate on-the-fly training convergence. However, previous methods, such as 3DGStream [38], typically avoid reusing these Gaussians to prevent the accumulation of an excessive number of Gaussians, which can lead to prohibitive training and streaming overhead. To address this limitation, we propose a selective inheritance mechanism for the densified Gaussian primitives $\bm{\mathcal{G}}^{\textrm{densi}}_{t-1}$. This approach adaptively selects a subset of the Gaussians that are beneficial for the reconstruction while controlling the total number of Gaussians to avoid excessive accumulation.
Specifically, for each Gaussian primitive in $\bm{\mathcal{G}}^{\textrm{densi}}_{t-1}$, we assign a learnable parameter. The parameter vector for these Gaussians is denoted by $\mathbf{m}$. We then apply the sigmoid function to $\mathbf{m}$, yielding $\textrm{sigmoid}(\mathbf{m})$, followed by quantization to generate a binary mask. This mask indicates the existence of inherited Gaussians, which is element-wise multiplied with the Gaussian opacities and scales as follows:
@@
where $\mathbf{o}_{r}$ and $\mathbf{s}_{r}$ are the opacity and scale values used during rendering, $\mathbf{o}$ and $\mathbf{s}$ are the original opacity and scale values, and $\circ$ represents the element-wise multiplication. When the mask is quantized to zero, the corresponding Gaussian has zero opacity and scale, thus contributing nothing to the rendering process. Therefore, this mask determines whether each Gaussian is retained in the rendering. By optimizing $\mathbf{m}$, the method selectively inherits the most relevant densified Gaussians from $\bm{\mathcal{G}}^{\textrm{densi}}_{t-1}$. The loss function for optimizing $\mathbf{m}$ is expressed as follows:
@@
The first two terms represent the fidelity loss in the vanilla 3DGS [21], where $\mathcal{L}_{1}$ and $\mathcal{L}_{\textrm{D-SSIM}}$ are pixel-level $L_{1}$ loss and D-SSIM [46] loss, respectively. The final term, weighted by $\lambda_{\textrm{inher}}$, is referred to as the mask loss and serves as a regularizer that encourages the parameters in $\mathbf{m}$ to approach zero. This regularization effectively reduces the number of inherited Gaussians and controls the numerical accumulation, while the fidelity loss aims to keep the Gaussians that are beneficial for reconstruction. Consequently, this trade-off optimizes the learnable $\mathbf{m}$ and selectively inherits important Gaussians from $\bm{\mathcal{G}}^{\textrm{densi}}_{t-1}$. After $s_{1}$ optimization steps, the final optimized mask $\textrm{Quant}(\textrm{sigmoid}(\mathbf{m}))$ is employed to remove redundant densified Gaussians and preserve the important ones as $\bm{\mathcal{G}}^{\textrm{inher}}_{t-1}$, which is subsequently fed into the following shift and densification stages.
This selective inheritance enhances the optimization pipeline in two key aspects. First, when training at timestep $t$ with the previous timestep’s base Gaussians $\bm{\mathcal{G}}^{\textrm{base}}_{t-1}$, distortions may occur due to temporal transformations and emerging objects. Feeding the base Gaussians directly into the shift stage can lead to inaccurate adjustments in Gaussian properties. Selectively inheriting the previous densified Gaussians $\bm{\mathcal{G}}^{\textrm{inher}}_{t-1}$ before the shift stage mitigates these optimization errors. Second, after the selective inheritance and shift stages, we obtain a refined set of optimized Gaussians $\bm{\mathcal{G}}^{\textrm{inher}}_{t}$, which reduces the number of densified Gaussians requiring optimization, thus improving training efficiency.
Figure 3: Histogram of Gaussian deformations in the flame steak scene of the N3DV dataset [26]. The overall distribution of Gaussian deformations (Yellow) is widely spread, with the majority falling into the low deformation range (less than 0.01). The dynamic (Orange) and static (Blue) components display different deformation patterns, where significant transformations are mainly concentrated in the dynamic component and minimal transformations are primarily found in the static component.
@@ 3.3 Dynamics-Aware Shift
In this stage, we focus on modeling the movements and rotations of Gaussian primitives from the preceding frame to the current one. A common approach to achieve this is directly learning deformation fields to accommodate temporal transformations, as in previous deformation-based method [47] and iterative streaming method [38]. However, these approaches often overlook the significant diversity in the movements of Gaussian primitives present in natural scenes, which can result in slower convergence. Specifically, Gaussians representing background or stationary object typically show high similarity with minimal variations or remain unchanged, while those in the foreground, such as humans, animals, and other moving objects, display significant dynamics. This disparity is illustrated in Fig. 3, which presents a histogram of Gaussian positional offsets between frames in the flame steak scene of N3DV [26]. We observe that most offsets fall within the range of $(0,0.01]$, which indicates that the majority of Gaussians undergo only minor shifts. Therefore, it is inappropriate to optimize the diverse variations in the scene using a single deformation field.
@@
To address this, we propose estimating a per-Gaussian dynamics mask for $\bm{\mathcal{G}}_{t-1}=\{\bm{\mathcal{G}}^{\textrm{base}}_{t-1},\bm{\mathcal{G}}^{\textrm{inher}}_{t-1}\}$ before deformation. This mask is used to categorize all primitives into dynamic and static groups. Then, we assign two deformation layers to learn the respective Gaussian transformations: A complex hash-encoding MLP [34] to capture the intricate shifts and rotations in the dynamic group, and a simpler hash-encoding MLP to model the regionally similar minor variations in the static group.
@@
To construct the per-Gaussian dynamics mask, we leverage established techniques from optical flow [48, 49, 45] and segmentation [23, 20, 53, 36]. Specifically, we utilize Gaussian Grouping [53] to initialize the scene $\bm{\mathcal{G}}_{0}$, which assigns an object property, akin to color property, to each Gaussian primitive and optimizes this property with 2D segmentation results serving as ground truth. These object properties provide per-Gaussian segmentation results that connect 2D images with 3D Gaussians, which facilitates the rendering of segmentation results on any viewpoint. Notably, these object properties remain fixed in subsequent frames $\bm{\mathcal{G}}_{t}(t>0)$, thus posing no additional burden on the time efficiency of on-the-fly training. For subsequent frames, we infer an optical flow estimation network [49] from consecutive frames captured from the same viewpoint and identify the areas where the optical flow exceeds a threshold $\gamma_{op}$, which we classify as dynamic areas. We then query the rendered segmentation results to find out the dynamic object IDs. The dynamic Gaussians are consistently detected using these IDs, as shown in Fig. 4. This process utilizes off-the-shelf methods that do not require network training, allowing us to obtain the per-Gaussian dynamics mask within milliseconds (about 300 ms). An introduction of these methods is provided in the supplementary material.
@@
After obtaining the dynamics mask, we employ the multi-resolution hash-encoding layer I-NGP [34] to learn the deformations. For each Gaussian, the deformation includes a position offset $\bm{\mu}_{n}\in\mathbb{R}^{3}$, which is added to the Gaussian position $\bm{p}_{n}$, and a rotation offset $\bm{\sigma}_{n}\in\mathbb{R}^{4}$, which is applied to the Gaussian rotation quaternion $\bm{q}_{n}$ by using $\mathrm{norm}({\bm{q}_{n}})\times\mathrm{norm}(\bm{\sigma}_{n})$. Here, $\mathrm{norm}(\cdot)$ denotes the normalization operation. The expressive ability of I-NGP is affected by the network complexity parameters, including the hash table size $T_{\textrm{Hash}}$ and the number of feature dimensions per entry $F_{\textrm{Hash}}$. For the dynamic group, we utilize a large-sized hash-encoding layer $\mathcal{H}_{\textrm{dyn}}$ with $T_{\textrm{Hash}}=T_{\textrm{large}}$ and $F_{\textrm{Hash}}=F_{\textrm{large}}$ to accommodate the complex variations of dynamic Gaussians. In contrast, for the static group, we employ a simpler hash-encoding layer $\mathcal{H}_{\textrm{st}}$ with $T_{\textrm{Hash}}=T_{\textrm{large}}/4$ and $F_{\textrm{Hash}}=F_{\textrm{large}}/2$, conserving computational resources while effectively managing minor jitters in background and stationary objects. These lightweight networks are optimized and provide a more efficient alternative to directly tuning the parameters of all Gaussians. To train $\mathcal{H}_{\textrm{dyn}}$ and $\mathcal{H}_{\textrm{st}}$, we use the fidelity loss function as follows:
@@ 3.4 Error-Guided Densification
The above two stages inherit temporally continuous Gaussians and manage deformations, without introducing new Gaussians. However, similar to the vanilla 3DGS [21] and 4DGS [52], densification remains essential to address areas with insufficient reconstruction quality and new objects.
@@
To achieve fast and high-quality densification, it is crucial to identify the subset of Gaussians in need of densification. Previous methods [21, 38] detect these under-reconstructed Gaussians based on historical view-space positional gradients, denoted by $\overline{\nabla\mathbf{p}}\triangleq\{\overline{\nabla p_{n}},\forall n\}\in\mathbb{R}^{N\times 1}$. This approach tracks gradients from previous training steps and densifies Gaussians whose positional gradients exceed a specific threshold $\tau_{\textrm{pos}}$. While this criterion effectively identifies general regions with inadequate reconstruction, it does not provide sufficient emphasis on regions with emerging objects or significant reconstruction errors. This limitation can hinder the efficient optimization of emerging objects, potentially slowing the convergence of on-the-fly training. To address this, we propose an error-adaptive densification strategy that incorporates an additional indicator to identify Gaussians that require densification. This approach projects high-distortion areas from the image space to the 3D space and adaptively enhances densification in critical areas.
@@
Specifically, we collect 2D error maps by comparing the ground truth with the rendered results in the shift stage for each training viewpoint, denoted by $\mathbf{E}^{c}$, $c=1,2,\cdots,C$. Then, we filter out the pixel positions with severe distortion above $\gamma_{\textrm{err}}$, yielding an $H\times W$ binary matrix $\mathbf{D}_{\textrm{err}}^{c}=\{\mathbf{1}[e_{i}^{c}>\gamma_{\textrm{err}}],\forall e^{c}_{i}\in\mathbf{E}^{c},i=1,\cdots,H\times W\}$ in the 2D pixel plane, where $\mathbf{1}[\cdot]$ is the binary indicator function. Next, we project the positions of the 3D Gaussians $\bm{\mathcal{G}}^{\textrm{base}}$ onto the 2D image plane as $(\mathbf{X},\mathbf{Y})=\text{Proj}(\bm{\mathcal{G}}^{\textrm{base}})$, where $\mathbf{X}\triangleq\{\mathbf{x}_{n},\forall n\},\mathbf{Y}\triangleq\{\mathbf{y}_{n},\forall n\}$, with $(\mathbf{x}_{n},\mathbf{y}_{n})$ representing the 2D pixel coordinate of the $n$-th Gaussian primitive along the $H$ and $W$ axes. The projection $\text{Proj}(\cdot)$ utilizes the known intrinsic and extrinsic camera matrices, with further details provided in the supplementary material. We then identify the subset of Gaussians that fall within the erroneous areas for each viewpoint, represented as $\mathcal{S}^{c}_{\textrm{err}}=\{\bm{\mathcal{G}}^{\textrm{base}}|\mathbf{D}_{\textrm{err}}^{c}(\mathbf{x}_{n},\mathbf{y}_{n})=1,\forall n\}$.
@@
For these highlighted Gaussian primitives, we apply a relatively lower gradient threshold $\tau_{\textrm{err}}<\tau_{\textrm{pos}}$, thereby placing greater emphasis on high-distortion areas and facilitating more effective error rectification compared to strategies that rely only on a single indicator. This leads to an error-guided adaptive densification scheme, yielding the subset of Gaussians for densification $\mathcal{S}$ as follows:
@@
where $\mathcal{S}_{\textrm{err}}=\{\cup^{C}_{c=1}\mathcal{S}^{c}_{\textrm{err}}\}$ is the combination of high-distortion subsets over all viewpoints. This error-guided adaptive subset $\mathcal{S}$ prioritizes the defects in historical optimization steps while accommodating under-reconstructed areas. We then perform spawn densification to compensate for both weakly reconstructed regions and emerging objects, following the vanilla 3DGS [21] and 3DGStream [38]. The newly added Gaussians, along with the transformed inherited Gaussians $\bm{\mathcal{G}}^{\textrm{inher}}_{t}$, are optimized using the fideltiy loss function as in Eq. (3) to yield $\bm{\mathcal{G}}^{\textrm{densi}}_{t}$. During this optimization, the pruning strategy is employed to eliminate excessive Gaussian candidates with very low opacity. Notably, these densified Gaussians $\bm{\mathcal{G}}^{\textrm{densi}}_{t}$ constitute only a small fraction of the total number of primitives in the scene, significantly saving computational and time costs compared to full-scene optimization.
3.5 Summary
Our proposed pipeline, DASS, effectively achieves iterative streamable 4D reconstruction through the seamless integration of the above three stages. The inheritance stage selectively preserves the optimized densified Gaussians from the previous frame. By exploiting the temporal consistency, this stage compensates for distortions caused by emerging objects before the shift stage and alleviates the optimization burden in the subsequent densification stage. The dynamics-aware shift stage incorporates separate deformation fields for dynamic and static features. This design is tailored to accommodate complex movements in the dynamic elements while conversing optimization resources for static components. After solving the temporal deformations, the error-guided densification stage reconstructs emerging objects in the current frame. It employs both positional gradients and rendering distortions as references to effectively identify the areas with weak reconstruction and emerging elements. Collectively, these three stages enable DASS to achieve high-fidelity reconstruction while significantly accelerating on-the-fly optimization.
4 Experiments
4.1 Experimental Setup
Implementation. We initiate our training at timestep 0 using Gaussian Grouping [53] and optimize the object properties for segmentation concurrently during this initialization, after which these properties remain fixed for subsequent frames. The optical flow is directly inferred using Unimatch [49]. These off-the-shelf methods obtain the segmentation and optical flow results without requiring extra pre-training, thereby imposing minimal computational or temporal overhead during on-the-fly training and streaming. For subsequent timesteps, the optimization processes for the inheritance, shift, and densification stages are executed with $s_{1}=20$, $s_{2}=100$, and $s_{3}=60$ steps, respectively.
@@ 4.1 Experimental Setup
Online StreamRF [25] 30.68 - 15 s 8.3 $\checkmark$
3DGStream [38] 31.67 - 12 s 215 $\checkmark$
DASS (Ours) 31.99 0.0285 9.62 s 189.2 $\checkmark$

### v2

@@ 3.2 Selective Inheritance
Exploiting temporal correlations is critical for effective and efficient Gaussian Splatting streaming. Due to the strong temporal consistency inherent in the scene, the base Gaussians $\bm{\mathcal{G}}^{\textrm{base}}$ and a subset of the temporally-varied Gaussians $\bm{\mathcal{G}}^{\textrm{var}}$ from the previous timestep remain applicable for subsequent timesteps, since they have been specifically optimized for the scene. Leveraging these timestep-wise priors and dependencies can significantly accelerate on-the-fly training convergence. However, previous methods, such as 3DGStream [37], typically retain the base Gaussians $\bm{\mathcal{G}}_{t-1}^{\textrm{base}}$ while discarding the previously densified Gaussians $\bm{\mathcal{G}}_{t-1}^{\textrm{var}}$ to prevent excessive accumulation of Gaussians, which can lead to prohibitive training and streaming overhead. To address this limitation, we propose a selective inheritance mechanism for $\bm{\mathcal{G}}_{t-1}^{\textrm{var}}$. Our approach adaptively selects a subset of the Gaussians that are beneficial for the current reconstruction while controlling the total number of Gaussians to avoid excessive accumulation.
Specifically, for each Gaussian primitive in $\bm{\mathcal{G}}^{\textrm{var}}_{t-1}$, we assign a learnable parameter, resulting in a parameter vector $\mathbf{m}\in\mathbb{R}^{N\times 1}$. We then apply the sigmoid function to $\mathbf{m}$, yielding $\textrm{sigmoid}(\mathbf{m})$, followed by quantization to generate a binary mask. This mask indicates the selection of inherited Gaussians and is element-wise multiplied with the Gaussian opacities and scales as follows:
@@
where $\mathbf{o}_{r}$ and $\mathbf{s}_{r}$ are the opacity and scale values used during rendering, $\mathbf{o}$ and $\mathbf{s}$ are the original opacity and scale values, and $\circ$ represents the element-wise multiplication. When the mask is quantized to zero, the corresponding Gaussian has zero opacity and scale, thus contributing nothing to the rendering process. Therefore, this mask determines whether each Gaussian is retained for rendering. By optimizing $\mathbf{m}$, the method selectively inherits the most relevant densified Gaussians from $\bm{\mathcal{G}}^{\textrm{var}}_{t-1}$. The loss function for optimizing $\mathbf{m}$ is expressed as follows:
@@
The first two terms represent the fidelity loss in the vanilla 3DGS [21], comprising pixel-level $L_{1}$ loss and D-SSIM [45] loss, respectively. The final term, weighted by $\lambda_{\textrm{inher}}$, is referred to as the mask loss and acts as a regularizer that encourages the parameters in $\mathbf{m}$ to approach zero. This regularization effectively reduces the number of inherited Gaussians and controls numerical accumulation, while the fidelity loss aims to retain Gaussians that are beneficial for reconstruction. Consequently, with only a few optimization iterations, this trade-off effectively optimizes the learnable $\mathbf{m}$ and selectively inherits important Gaussians from $\bm{\mathcal{G}}^{\textrm{var}}_{t-1}$. The final optimized mask $\textrm{Quant}(\textrm{sigmoid}(\mathbf{m}))$ is employed to remove redundant densified Gaussians and preserve the important ones, which are then fed into the shift stage.
This selective inheritance mechanism enhances the optimization pipeline in two key aspects. First, the differences between timesteps arise from both temporal transformations and emerging objects. Directly feeding only the base Gaussians into shift stage can lead to inaccurate adjustments in Gaussian properties. By selectively inheriting the previous densified Gaussians before the shift stage, we mitigate these optimization errors. Second, after the selective inheritance and shift stages, we obtain a refined set of optimized Gaussians, which reduces the number of densified Gaussians requiring optimization, thus improving training efficiency.
Figure 3: Histogram of Gaussian deformations in the flame steak scene of the N3DV dataset. The overall distribution of Gaussian deformations (Yellow) is widely spread, with the majority falling into the low deformation range (less than 0.01). The dynamic (Orange) and static (Blue) components display different deformation patterns, where significant transformations are mainly concentrated in the dynamic component and minimal transformations are primarily found in the static component.
@@ 3.3 Dynamics-Aware Shift
In this stage, we focus on modeling the movements and rotations of Gaussian primitives from the preceding timestep to the current one. A common approach to achieve this is directly learning deformation fields to accommodate temporal transformations, as in previous deformation-based method [46] and iterative streaming method [37]. However, these approaches often overlook the significant diversity in the movements of Gaussians present in natural scenes, which can result in slower convergence. Specifically, Gaussians representing background or stationary object typically show high similarity with minimal variations or remain unchanged, while those in the foreground, such as humans and other moving objects, display significant dynamics. This disparity is illustrated in Fig. 3, presenting a histogram of Gaussian positional offsets between timesteps in the flame steak scene of N3DV [26]. We observe that most offsets fall within the range of $(0,0.01]$, which indicates that the majority of Gaussians undergo only minor shifts. Therefore, it is inappropriate to optimize the diverse variations in the scene using a single deformation field.
@@
To address this, we propose estimating a per-Gaussian dynamics mask for all Gaussians before deformation. This mask is used to categorize all primitives into dynamic and static groups. Then, we assign two deformation layers to learn the respective Gaussian transformations: A complex hash-encoding MLP [33] to capture the intricate shifts and rotations in the dynamic group, and a simpler hash-encoding MLP to model the regionally-similar minor variations in the static group.
@@
To construct the per-Gaussian dynamics mask before deformation, we leverage techniques from optical flow [47, 48, 44] and segmentation [23, 20, 52, 35]. Specifically, we utilize Gaussian Grouping [52] to initialize the scene $\bm{\mathcal{G}}_{0}$, which assigns an object property, akin to color property, to each Gaussian primitive and optimizes this property with 2D segmentation results as ground truth. These object properties provide per-Gaussian segmentation results that connect 2D images with 3D Gaussians, which facilitates the rendering of segmentation results on any viewpoint. Notably, these object properties remain fixed in subsequent timesteps $\bm{\mathcal{G}}_{t}(t>0)$, thus posing no additional burden on the time efficiency of on-the-fly training. For subsequent timesteps, we infer an optical flow estimation network [48] on consecutive timestep frames captured from the same viewpoint and identify the areas where the optical flow exceeds a threshold $\gamma_{op}$, which we classify as dynamic areas. We then query the rendered segmentation results to find out the dynamic object IDs. The dynamic Gaussians are consistently detected using these IDs, as shown in Fig. 4. This process utilizes off-the-shelf methods that do not require network training, allowing us to obtain per-Gaussian dynamics mask within milliseconds (about 300 ms).
@@
After obtaining the dynamics mask, we employ the multi-resolution hash-encoding layer I-NGP [33] to learn the deformations. For each Gaussian, the deformation includes a position offset $\bm{\mu}_{n}\in\mathbb{R}^{3}$, which is added to the Gaussian position $\bm{p}_{n}$, and a rotation offset $\bm{\sigma}_{n}\in\mathbb{R}^{4}$, which is applied to the Gaussian rotation quaternion $\bm{q}_{n}$ by using $\mathrm{norm}({\bm{q}_{n}})\times\mathrm{norm}(\bm{\sigma}_{n})$. Here, $\mathrm{norm}(\cdot)$ denotes the normalization operation. To handle the distinct characteristics of dynamic and static Gaussians, we employ a dual-network strategy: a high-capacity hash-encoding layer $\mathcal{H}_{\textrm{dyn}}$ is used for the dynamic group to capture their complex and varied deformations, while a simpler, lightweight hash-encoding layer $\mathcal{H}_{\textrm{st}}$ is applied to the static group to efficiently manage minor jitters in background and stationary objects. This approach not only conserves computational resources but also provides an effective alternative to directly tuning the parameters of all Gaussians. The networks $\mathcal{H}_{\textrm{dyn}}$ and $\mathcal{H}_{\textrm{st}}$ are trained using the fidelity loss function as follows:
@@
Figure 5: (Left) Gaussians identified for densification based on positional gradients. (Right) Gaussians identified for densification using error-guidance distortion projection. This comparison verifies that our proposed strategy effectively prioritizes emerging objects and achieves targeted compensations. While the vanilla densification strategy requires multiple optimization steps to fully recover the scene, our method concentrates on high-distortion areas with emerging objects, enhancing the computational efficiency.
@@ 3.4 Error-Guided Densification
The above two stages focus on maintaining temporal continuity and managing deformations without introducing new Gaussians. Besides these two stages, densification remains essential to address areas with insufficient reconstruction quality and to incorporate new objects.
@@
To achieve fast and high-quality densification, it is crucial to identify the subset of Gaussians in need of densification. Previous methods [21, 37] detect under-reconstructed Gaussians based on historical view-space positional gradients, denoted by $\overline{\nabla\mathbf{p}}\triangleq\{\overline{\nabla p_{n}},\forall n\}\in\mathbb{R}^{N\times 1}$. This approach tracks gradients from previous training steps and densifies Gaussians whose positional gradients exceed a specific threshold $\tau_{\textrm{pos}}$. While this criterion effectively identifies general regions with inadequate reconstruction, as shown in Fig. 5 (Left), it fails to provide sufficient emphasis on regions with emerging objects or significant reconstruction errors. This limitation can hinder the efficient optimization of emerging objects, potentially slowing the convergence of on-the-fly training. To address this, we propose an error-adaptive densification strategy that incorporates an additional indicator to identify Gaussians requiring densification. This approach projects high-distortion areas from the image space to the 3D space and adaptively enhances densification in critical areas.
@@
Specifically, we collect 2D error maps by comparing the ground truth with the rendered results in the shift stage for each training viewpoint, denoted by $\mathbf{E}^{c}$, $c=1,2,\cdots,C$. Then, we filter out the areas with severe distortion above $\gamma_{\textrm{err}}$, yielding an $H\times W$ binary matrix $\mathbf{D}_{\textrm{err}}^{c}=\{\mathbf{1}[e_{i}^{c}>\gamma_{\textrm{err}}],\forall e^{c}_{i}\in\mathbf{E}^{c},i=1,\cdots,H\times W\}$ in the 2D pixel plane, where $\mathbf{1}[\cdot]$ is the binary indicator function. Next, we project the positions of the 3D Gaussians onto the 2D image plane as $(\mathbf{X},\mathbf{Y})=\text{Proj}(\bm{\mathcal{G}}^{\textrm{base}})$, where $\mathbf{X}\triangleq\{\mathbf{x}_{n},\forall n\},\mathbf{Y}\triangleq\{\mathbf{y}_{n},\forall n\}$, with $(\mathbf{x}_{n},\mathbf{y}_{n})$ representing the 2D pixel coordinate of the $n$-th Gaussian primitive along the $H$ and $W$ axes. The projection $\text{Proj}(\cdot)$ utilizes the known intrinsic and extrinsic camera matrices, with further details provided in the supplementary material. We then identify the subset of Gaussians that fall within the erroneous areas for each viewpoint, represented as $\mathcal{S}^{c}_{\textrm{err}}=\{\bm{\mathcal{G}}^{\textrm{base}}|\mathbf{D}_{\textrm{err}}^{c}(\mathbf{x}_{n},\mathbf{y}_{n})=1,\forall n\}$.
@@
For these highlighted Gaussian primitives, we apply a relatively lower gradient threshold $\tau_{\textrm{err}}<\tau_{\textrm{pos}}$, thereby placing greater emphasis on high-distortion areas and facilitating more effective error rectification compared to strategies that rely only on positional gradients. This leads to an error-guided adaptive densification scheme, yielding the subset of Gaussians for densification $\mathcal{S}$ as follows:
@@
where $\mathcal{S}_{\textrm{err}}=\{\cup^{C}_{c=1}\mathcal{S}^{c}_{\textrm{err}}\}$ is the combination of high-distortion subsets over all viewpoints (Fig. 5 (Right)). This error-guided adaptive subset $\mathcal{S}$ prioritizes the defects in historical optimization steps while accommodating under-reconstructed areas. We then perform spawn densification to compensate for both weakly reconstructed regions and emerging objects. The newly added Gaussians and the transformed inherited Gaussians from $\bm{\mathcal{G}}^{\textrm{var}}_{t-1}$ are optimized using the fidelity loss function in Eq. (3) to yield $\bm{\mathcal{G}}^{\textrm{var}}_{t}$. During this optimization, the pruning strategy is employed to eliminate excessive Gaussian candidates with very low opacity. Notably, these temporally-varied Gaussians constitute only a small fraction of the total number of primitives in the scene, significantly reducing computational and time costs compared to full-scene optimization.
@@
Online StreamRF [25] 30.71 - 14.20 s 9 $\checkmark$
3DGStream [37] 31.68 0.0299 12.41 s 215 $\checkmark$
HiCoM [15] 31.30 0.0316 10.22 s 230 $\checkmark$
DASS (Ours) 31.99 0.0285 9.62 s 207 $\checkmark$
Figure 6: (Left) Qualitative comparison for the scene coffee martini of N3DV dataset, including online training methods [25, 37] and offline reconstruction method [46]. Compared to other methods, our method accurately recovers dynamic objects and preserves intricate details. (Right) Error map visualization without (top) and with (bottom) the proposed selective inheritance mechanism, on scene flame steak. The selective inheritance effectively compensates the artifacts in temporally consistent areas and mitigates the optimization distortions.
4 Experiments
4.1 Experimental Setup
Implementation. We initiate our training at timestep 0 using Gaussian Grouping [52] and optimize the object properties for segmentation concurrently during this initialization, after which these properties remain fixed for subsequent timesteps. The optical flow is directly inferred using Unimatch [48]. These off-the-shelf methods obtain the segmentation and optical flow results without extra pre-training, thereby imposing minimal computational or temporal overhead during on-the-fly training and streaming. For subsequent timesteps, the optimization processes for the inheritance, shift, and densification stages are executed with $s_{1}=20$, $s_{2}=100$, and $s_{3}=60$ steps, respectively.